OpenStudy (anonymous):

simplify the radical expression sqareroot 180x^2

OpenStudy (solomonzelman):

\[180=36 \times 5\]

OpenStudy (anonymous):

The Answer is: \[6x^2\sqrt{5}\]

OpenStudy (anonymous):

6sqrt5x^2?

OpenStudy (solomonzelman):

No it is not

OpenStudy (anonymous):

oh ok thank you

OpenStudy (solomonzelman):

Incorrect.

OpenStudy (anonymous):

\[6x^2*\sqrt{5}\] i Mean

OpenStudy (anonymous):

what is it solomon

OpenStudy (solomonzelman):

no

OpenStudy (anonymous):

was what i posted correct?

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \sqrt{180x^2}=\sqrt{5 \times36x^2} =\sqrt{5}\times \sqrt{(6x)^2} =6x\sqrt{5} }\)

OpenStudy (anonymous):

Ok Was The Question Simplify : \[\sqrt{180x^4}\]?

OpenStudy (anonymous):

is it 6xsqrt5?

OpenStudy (anonymous):

Solomon Thats What I Said -.-

OpenStudy (anonymous):

The Answer is : \[6x \sqrt{5}\]